OpenStudy (anonymous):

Factor by grouping to obtain the difference of two squares. Then, factor completely. x^2-18x+81-y^2

OpenStudy (zarkon):

\[x^2-18x+81=(x-9)^2\] then \[x^2-18x+81-y^2=(x-9)^2-y^2=(x-9+y)(x-9-y)\]